OpenStudy (anonymous):

The sequence –5, 0, 5, 10,..., 65 has 15 terms. Evaluate the related series. Is it 450?

Parth (parthkohli):

Series is the sum of all the terms.

OpenStudy (anonymous):

I got 450 as my answer, i that correct?

Parth (parthkohli):

10 + 15 + 20 + 25 + 30 + 35 + 40 + 45 + 50 + 55 + 60 + 65

Parth (parthkohli):

Yes. it is 450

Parth (parthkohli):

You can also use sigma notation to solve.

OpenStudy (anonymous):

Sn = (a1+an)*n/2 Sn = (-5 + 65) *15 / 2 = 450